# Targeted Fine-Tuning of DNN-Based Receivers via Influence Functions

### v1

@@ 2 Background
Influence Functions (IFs) originate from robust statistics [22] and were adapted to deep learning by Koh and Liang [21]. They estimate how perturbing a training instance $z_{i}$, $i\in\{1,\ldots,n\}$, changes the learned parameters and, in turn, the loss on a test point $z_{\text{test}}$. For empirical risk minimization with loss $\ell$, let $\hat{\theta}$ denote the learned parameters. The parameter change from upweighting $z_{i}$ by an infinitesimal amount $\epsilon$ is [21]
@@ 2.1 Cross-Loss Influence – Aligning with Task Metrics
Cross-loss IF (CLIF) [26] generalizes the framework by decoupling the training and evaluation losses. For training loss $\ell_{\text{train}}$ and evaluation loss $\ell_{\text{eval}}$, the influence is [26]
@@ 2.2 Relative Influence – Highlighting Local Influence
Relative influence functions [27] constrain the comparison either in parameter space or in loss space:
@@ 2.4 Influence Approximations – Scaling to Deep Models
Computing IFs requires Inverse-Hessian-Vector Products (IHVPs), which are intractable for large networks. Arnoldi-Based IF [25] projects the Hessian onto a low-dimensional Krylov subspace:
$H_{\hat{\theta}}^{-1}v\approx QT^{-1}Q^{\top}v=\sum_{i=1}^{k}\tfrac{1}{\lambda_{i}}u_{i}(u_{i}^{\top}v),$ (7)
where $Q$ is an orthonormal basis, $T$ a small Hessenberg matrix from the Arnoldi process [31] with eigenpairs $(\lambda_{i},y_{i})$, and $u_{i}=Qy_{i}$. This yields efficient Ritz approximations [31] of $H_{\hat{\theta}}^{-1}v$ for IHVPs.
@@ 3 Targeted fine-tuning
Fig. 2: Relative BER gaps $\Delta_{\text{BER}}$ as defined in Eq. (9) between DeepRx and a genie-aided benchmark. Targets are the top-5 largest-gap instances in the regime where $\text{BER}_{\text{Genie}}\geq 10^{-3}$.
@@ 5 Discussion
In multi-target settings, influence-guided tuning offered limited benefit, often matching but not surpassing random tuning. Group-based approaches [37] may provide more coherent updates across multiple targets. Beyond this, advanced fine-tuning methods such as LoRA [38] with PEFT [39] or trajectory-based influence estimators like TracIn [40] are also promising directions for future research.
Beyond fine-tuning, influence functions enhance interpretability by revealing which training subsets or input features (e.g., subcarriers) most affect predictions [35, 41]. They also suggest new training strategies: receivers could first be trained on smaller subsets and then refined with influential samples to reduce data costs, or adapted on the fly to new channel conditions without full retraining—supporting more robust and efficient deployment in dynamic wireless environments.
@@ 6 Conclusions
Open questions remain on (i) the choice of influence loss, (ii) criteria for selecting poorly performing instances, (iii) how to exploit harmful instances in fine-tuning, and (iv) how to handle multiple targets jointly. Promising future directions include lightweight adaptation methods (e.g., LoRA with PEFT), alternative influence estimators (e.g., TracIn), and group-based selection strategies.

### v2

@@ 2 Background
Influence Functions (IFs) originate from robust statistics [22] and were adapted to deep learning by Koh and Liang [21]. They estimate how perturbing a training instance $z_{i}$, $i\in\{1,\ldots,n\}$, changes the learned parameters and, in turn, the loss on a test point $z_{\text{test}}$. For empirical risk minimization with loss $\ell$, let $\hat{\theta}$ denote the learned parameters. The parameter change from upweighting $z_{i}$ by an infinitesimal amount $\epsilon$ is [21, Eq.1]
@@ 2.1 Cross-Loss Influence – Aligning with Task Metrics
Cross-loss IF (CLIF) [26] generalizes the framework by decoupling the training and evaluation losses. For training loss $\ell_{\text{train}}$ and evaluation loss $\ell_{\text{eval}}$, the influence is [26, Theorem 1]
@@ 2.2 Relative Influence – Highlighting Local Influence
Relative influence functions constrain the comparison either in parameter space or in loss space [27, Definitions 3.2 and 3.4]:
@@ 2.4 Influence Approximations – Scaling to Deep Models
Computing IFs requires Inverse-Hessian-Vector Products (IHVPs), which are intractable for large networks. Arnoldi-Based IF [25] projects the Hessian onto a low-dimensional Krylov subspace [31]:
$H_{\hat{\theta}}^{-1}v\approx G^{\top}\widetilde{H}^{-1}Gv=\sum_{i=1}^{k}\tfrac{1}{\lambda_{i}}g_{i}(g_{i}^{\top}v),$ (7)
where $G$ is a projection, $\widetilde{H}=GH_{\hat{\theta}}G^{\top}$ is projected Hessian induced by Arnoldi process [31] with $k$ eigenpairs $(\lambda_{i},y_{i})$, and $g_{i}=G^{\top}y_{i}$. This yields efficient Ritz approximations [31] for IHVPs.
@@ 3 Targeted fine-tuning
Fig. 2: Relative BER gaps $\Delta_{\text{BER}}$ as defined in Eq. (9) between DeepRx and a genie-aided LMMSE (full CSI). Targets are the top-5 largest-gap instances in the regime where $\text{BER}_{\text{Genie}}\geq 10^{-3}$.
@@ 5 Discussion
In multi-target settings, influence-guided tuning offered limited benefit, often matching but not surpassing random tuning. Group-based [37, Eq.4] and joint [20, Eq.3] approaches may yield more coherent multi-target updates. Beyond this, advanced fine-tuning such as LoRA [38] with PEFT [39] and trajectory-based influence estimators like TracIn [40] remain promising research directions.
Beyond fine-tuning, influence functions could enhance interpretability by revealing which training subsets or input features (e.g., subcarriers, time slots) most affect predictions [35, 41]–potentially linking patterns to communication-theoretic effects (e.g., frequency selectivity, pilot contamination). They also suggest data-efficient training (pretraining on smaller subsets, refining with influential samples) and on-the-fly adaptation–supporting more robust and efficient deployment in dynamic wireless environments.
@@ 6 Conclusions
Open questions remain on (i) the choice of influence loss, (ii) criteria for selecting poorly performing instances, (iii) how to exploit harmful instances in fine-tuning, and (iv) how to handle multiple targets jointly. Promising future directions include lightweight adaptation methods (e.g., LoRA with PEFT), alternative influence estimators (e.g., TracIn), and group- or joint-update strategies.